# Roman Garipov ${}^{*\,\dagger}$

### v1

@@ 3.3 Inference
The resulting classifier can be used with arbitrary speculative decoding algorithm that has a verification stage. During said verification stage, the classifier is called when the original algorithm would reject a token. If the would-be-rejected token is deemed to be unimportant, i.e. not to affect the response quality, then we override the verification procedure and accept the token instead, proceeding to test subsequent tokens (if any) as per the original algorithm.
Generality. In our initial experiments, we focus on traditional speculative decoding (Leviathan et al., 2023; Chen et al., 2023) for simplicity. However, our algorithm is compatible with arbitrary speculative decoding algorithms, including tree-based (Miao et al., 2023; Svirschevski et al., 2024; Chen et al., 2024) and single-model multi-head algorithms (Cai et al., 2024; Li et al., 2024b, a). This also means that our approach can be integrated into existing speculative decoding software such as vLLM (Kwon et al., 2023), TensorRT-LLM (NVIDIA, 2023) or TGI (Hugging Face, 2023).
Thresholds. To balance computational efficiency and downstream performance, we select a decision threshold that achieves a high recall (${\geq}90{\%}$) in order to retain quality. Since the classifier is accurate enough, this threshold can also achieve decent rejection rate, i.e., the rate of tokens correctly predicted to be unimportant. This allows us to retain downstream accuracy while skipping a large portion of unimportant tokens, thus enabling efficient speculative decoding. In Section 4, we also evaluate with various threshold values to show their effect on accuracy and acceptance rate.
Comparison with Judge Decoding. As we discussed earlier, our approach can be seen as an extension of Judge Decoding that enables automatic dataset mining. As such, the dataset generation algorithm from Section 3.1 can be used in conjunction with the Judge Decoding training and inference protocol, which appears to be similar to ours up to possible minor details. It would be interesting to compare the two dataset collection strategies directly, ceteris paribus.
@@ 4 Experiments
We evaluate the proposed approach in two setups: mathematical reasoning with GSM8K dataset (Cobbe et al., 2021) and programming with LiveCodeBench (Jain et al., 2024). In both cases, use the popular Llama 3.x model family, with Llama-3.1-8B-Instruct as the main model and Llama-3.2-1B-Instruct as the draft model [^4] The reason why the two models have different minor versions (i.e. 3.1 and 3.2) is that the 3.2 version does not have the larger 8B models and the 3.1 version does not have the smaller 1B models.. We run AutoJudge on top of standard speculative decoding algorithm Leviathan et al. (2023) with an extended draft size of 64 tokens. Our main experiments run in native bfloat16 precision, but we have found several peculiarities related to numeric precision, reported in Appendix C. We report GSM8K results in Section 4.1 and LiveCodeBench in Section 4.2.
@@ 4.1 Mathematical Reasoning with GSM8K
Our first set of experiments is based on the GSM8K dataset with grade school math problems. This dataset has a natural split with ${\approx}7.47$K training samples and ${\approx}1.32$K test samples. Following the standard evaluation procedure, we use the training set to “mine” important tokens with Algorithm 1 and train the classifier, then run inference and evaluate on the test set with the recommended parameters Gao et al. (2021) for zero-shot evaluation: greedy inference with a prompt that encourages chain-of-thought reasoning. During training, we consider two responses equivalent ($a{\equiv}\hat{a}$ in Alg. 1) if the extracted final answers (numbers) are equal. For reference, we provide an example important token assignments found by our algorithm in Figure 2.
@@
We train a classifier on last hidden state embeddings from both draft and target models (concatenated) for encoded draft tokens. The training dataset from ${\approx}7.47$K original samples contains ${\approx}130$K mismatches, about $20\%$ of which are deemed important. We train logistic regression with $C{=}10^{-4}$ regularization coefficient ($L_{2}$), found by grid search over a logarithmic grid between $10^{0}\dots 10^{-9}$.
During inference, we integrate the trained classifier into the speculative decoding loop from Leviathan et al. (2023) during verification. Whenever the original algorithm would reject a token, we run the classifier to determine if changing that token affects the final response quality, and if not — accept the token and continue verification for subsequent tokens (if any). Since the resulting algorithm can accept additional tokens, we use the increased draft window size of $W{=}64$ tokens for all evaluations. We report two main metrics: downstream accuracy and the number of accepted tokens per speculative decoding cycle. The accuracy is measured as the exact match rate for the final answer extracted from the response as per standard GSM8K evaluation protocol. In turn, we report decoding speed in terms of the number of tokens accepted per target model forward pass with the same speculative decoding parameters, so as to decouple our results from the specific hardware configuration.
@@
We evaluate AutoJudge with different classifier thresholds, balancing between accuracy and speed. Our baselines are traditional speculative decoding, decoding with the draft model and a simpler lossy speculative decoding protocol. In the latter, we accept a mismatching draft token if it is within top-$K$ most likely tokens of the target model, similarly to how it is defined in Bachmann et al. (2025). We report $K{=}2,4,8,\dots,|V|$ for different speed-accuracy trade-offs.
@@
The results in Figure 3 (left) demonstrate that AutoJudge decoding can achieve substantial speed-ups over both autoregressive inference and traditional speculative decoding. Varying the classifier threshold allows us to achieve both near-lossless accuracy with moderate speed-ups and even greater speed-ups at the cost of several percentage points drop in accuracy. The heuristic-based Top-$K$ baseline also achieves some speed-ups, but at the cost of significantly higher accuracy drawdown.
@@ 4.2 Programming with LiveCodeBench
Next, we test if AutoJudge search algorithm is able to generalize between domains. For this purpose, we evaluate the same model pair on LiveCodeBench Jain et al. (2024). For this evaluation, we use the code_generation_lite [^5] https://huggingface.co/datasets/livecodebench/code_generation_lite dataset with version tag release_v5. The dataset contains 880 programming tasks (we evaluate on all three subsets: easy, medium and hard). Since LiveCodeBench does not have a dedicated training split, we evaluate using out-of-fold predictions. Namely, we split the dataset randomly into 5 folds. For each fold, we evaluate using the classifier trained on the 4 remaining folds. We use the standard evaluation protocol: extracting the generated code and evaluating it using the benchmark’s builtin test suite.
[…]
@@
The results in Figure 3 (right) are similar to what we observed in Section 4.1: AutoJudge decoding can accept over 20 tokens per forward pass at the cost of ${\approx}1\%$ accuracy drawdown. This results in approximately 2×2$\times$2 × increase over traditional speculative decoding Leviathan et al. (2023). The Top-$K$ baseline can similarly achieve some increase in the number of accepted tokens, but AutoJudge decoding offers significantly better quality-speed trade-offs across all configurations. We report additional configurations and threshold values in Appendix D. We also evaluate AutoJudge decoding “out-of-domain”: using the classifier trained on GSM8K data for LiveCodeBench evaluation(also in Appendix D), which results in inferior performance. This aligns with our hypothesis that the important tokens depend on the problem type and evaluation criteria.
5 Discussion
@@
In this working paper, we propose and evaluate a fully automated protocol for task-specific speculative decoding acceleration. Our initial experiments suggest that a simple-based procedure can successfully determine which of the mismatching tokens in the LLM response affect the downstream quality for both mathematical reasoning and programming tasks. In the upcoming update, we plan to evaluate more practical speculative decoding setups by reporting results of experiments with larger target models. We hope that AutoJudge can facilitate the use of Judge Decoding across different tasks types, languages and modalities.
@@
In future work, we aim to explore the performance of AutoJudge decoding across additional tasks and models and analyze how its outputs differ depending on the use case. While it is, unfortunately, not possible to directly compare against the original Judge Decoding (see the end of Section 3), it would nonetheless be interesting to compare human and automated annotations in this setting. We also plan to explore how AutoJudge pairs with more advanced speculative decoding algorithms, such as speculative decoding with tree-based drafts (Miao et al., 2023; Chen et al., 2024; Svirschevski et al., 2024) or learned drafting heads (Cai et al., 2024; Li et al., 2024b, 2025). Finally, we plan to evaluate the runtime of speculative decoding with AutoJudge-based classifier in efficient frameworks such as vLLM (Kwon et al., 2023; NVIDIA, 2023; Hugging Face, 2023).
Appendix A Additional considerations from Section 3.1
@@
[…]
where GumbelPRNG is a function that samples a pseudo-random variable from standard Gumbel distribution based on an input seed $s\oplus x\oplus y_{1:t}$. To recall, $\oplus$ denotes concatenation. This way, $y_{next}$ is distributed as $P(y_{next}|x\oplus y_{1:t},\theta)$, but it is deterministic when conditioned on the random state $s$. Hence, we sample a random state $s$ once at the beginning of Algorithm1, the entire procedure after that will also be conditionally deterministic (given $s$).
@@ Appendix B Additional Details on Classifier Training
For the first set of experiments (Figure 4), we compare regularizer coefficients for Logistic Regression (left). We also report different classifier types: Logistic Regression, a Random Forest with 128 trees and a multi-layer perceptron (MLP) with a single hidden layer consisting of 128 hidden units with ReLU activation. For consistency, we run all models using Scikit-Learn Pedregosa et al. [2011] v1.4.2 with all other settings kept to their default values. For MLP, we perform early stopping on yet another 10% subset of the training set with built-in default MLPClassifier early stopping parameters. For this evaluation, all classifiers use draft and target model hidden states (concatenated) encoding the draft token, which is our main setup from Section 3.2.
@@ Appendix C Precision Matters for Speculative Decoding
Appendix D Additional Evaluations for Sections 4.1 & 4.2
In Figure 7, we report additional threshold configurations for AutoJudge and additional values of $K$ for the Top-$K$ baseline, extending Figure 3. Additionally, we evaluate the AutoJudge classifier trained on LiveCodeBench on GSM8K and vice versa to gauge the effect of task-specific training. Predictably, these out-of-domain classifiers perform significantly worse. We attribute this to the fact that the GSM8K-trained classifier likely did not see any Python source code, whereas the LiveCodeBench classifier did not perform arithmetic operations and did not solve equations that are common in GSM8K. In future, it would be interesting to explore combined classifier training (e.g. both math and code) to see if our approach is able to generalize to unseen tasks.

### v2

@@ 3.3 Inference
The resulting classifier can be used with an arbitrary speculative decoding algorithm that has a verification stage. During said verification stage, the classifier is called when the original algorithm would reject a token. If the would-be-rejected token is deemed to be unimportant, i.e. not to affect the response quality, then we override the verification procedure and accept the token instead, proceeding to test subsequent tokens (if any) as per the original algorithm.
Generality. In our initial experiments, we focus on traditional speculative decoding (Leviathan et al., 2023; Chen et al., 2023) for simplicity. However, our algorithm is compatible with arbitrary speculative decoding algorithms, including tree-based (Miao et al., 2023; Svirschevski et al., 2024; Chen et al., 2024) and single-model multi-head algorithms (Cai et al., 2024; Li et al., 2024b, a). This also means that our approach can be integrated into existing inference frameworks such as vLLM (Kwon et al., 2023), TensorRT-LLM (NVIDIA, 2023) or TGI (Hugging Face, 2023).
Thresholds. To balance computational efficiency and downstream performance, we select a decision threshold that achieves a high recall (${\geq}90{\%}$) in order to retain quality. Since the classifier is accurate enough, this threshold can also achieve a reasonable rejection rate, i.e., the rate of tokens correctly predicted to be unimportant. This allows us to retain downstream accuracy while skipping a large portion of unimportant tokens, thus enabling efficient speculative decoding. In Section 4, we evaluate various threshold values to show their effect on accuracy and acceptance rate.
Comparison with Judge Decoding. As we discussed earlier, our approach can be seen as an extension of Judge Decoding that enables automatic dataset mining. As such, the dataset generation algorithm from Section 3.1 can be used in conjunction with the Judge Decoding training and inference protocol, which appears to be similar to ours up to possible minor details. It would be interesting to compare the two dataset collection strategies directly. Unfortunately, the annotated dataset and the source code of Judge Decoding are not available, which makes such comparison difficult.
@@ 4 Experiments
Figure 3: Accuracy and the average number of accepted tokens on GSM8K for (left) 8-shot Llama-3.2 1B draft / Llama-3.1 8B target and (middle) 0-shot Llama 3.1 8B draft / Llama 3.1 70B target and (right) 8-shot Qwen2.5-0.5B draft / Qwen2.5-7.5B target. All models are Instruct versions.
We organize our evaluations as follows: in Section 4.1 we evaluate AutoJudge on the GSM8K (Cobbe et al., 2021) mathematical reasoning benchmark. Next, in Section 4.2, we evaluate our approach on programming tasks from LiveCodeBench (Jain et al., 2024). We use Llama 3.x model family across all setups, focusing on two model pairs: 1) Llama-3.2-1B-Instruct draft / Llama-3.1-8B-Instruct target∗*∗**The reason why the two models have different minor versions (i.e. 3.1 and 3.2) is that the 3.2 version does not have the larger 8B models, and the 3.1 version does not have the smaller 1B models. and 2) Llama-3.1-8B-Instruct draft / Llama-3.1-70B-Instruct target. Additionally, we report results with Qwen2.5-0.5B (draft) / Qwen2.5-7.5B (target) on the GSM8K benchmark to showcase the method’s transferability across model families. Finally, Section 4.3 contains GPU inference speed benchmarks with our vLLM implementation where we consider Llama-3.1-8B-Instruct draft / Llama-3.1-70B-Instruct target and Llama-3.1-8B-Instruct draft / Llama-3.1-405B-Instruct target setups. The 405B setup is omitted in Section 4.1 and Section 4.2, since those sections rely on transformers-based inference, which is significantly slower than vLLM. We run AutoJudge on top of the standard speculative decoding algorithm (Leviathan et al., 2023). Our main experiments run in bfloat16 precision (see Appendix C for details). We provide results for additional model pairs in Appendix D, extended inference benchmarks under various setups in Appendix F, evaluations with different window sizes in Appendix G, and inference with offloading experiments in Appendix H.
@@ 4.1 Mathematical Reasoning with GSM8K
Our first set of experiments is based on the GSM8K dataset with grade school mathematical problems. This dataset has a natural split with ${\approx}7.47$K training samples and ${\approx}1.32$K test samples. Following the standard evaluation procedure, we use the training set to “mine” important tokens with Algorithm 1 and train the classifier, then run inference and evaluate on the test set with the recommended parameters (Gao et al., 2021) for zero-shot and 8-shot evaluation: greedy inference with a prompt that encourages chain-of-thought reasoning. During training, we consider two responses equivalent ($a{\equiv}\hat{a}$ in Alg. 1) if the extracted final answers (numbers) are equal. For reference, we provide an example of important token assignments found by our algorithm in Figure 2.
@@
We train a classifier on the last hidden state embeddings from both draft and target models (concatenated) for encoded draft tokens. The training dataset contains ${\approx}130$K mismatches, about $20\%$ of which are deemed important. We train a logistic regression with the $L_{2}$ regularization coefficient tuned individually for each setup(for instance $10^{-4}$ for 8B/70B), found by grid search over a logarithmic grid between $10^{0},\dots,10^{-7}$. During inference, we integrate the trained classifier into the speculative decoding loop from Leviathan et al. (2023) during verification. Whenever the original algorithm would reject a token, we run the classifier to determine if changing that token affects the final response quality, and if not — accept the token and continue verification for subsequent tokens (if any). Since the resulting algorithm can accept additional tokens, we use the increased draft window size of W=64W{=}64 tokens for all evaluations.
@@
We report two main metrics: downstream accuracy and the number of accepted tokens per speculative decoding cycle. The accuracy is measured as the exact match rate for the final answer extracted from the response as per standard GSM8K evaluation protocol. In turn, we report the decoding speed in terms of the number of tokens accepted per target model forward pass, aiming to decouple our results from the specific hardware configuration. We evaluate AutoJudge with different classifier thresholds, balancing between accuracy and speed. Our baselines are traditional speculative decoding, decoding with the draft model, and a simpler lossy speculative decoding method. In the latter, we accept a mismatching draft token if it is within top-$K$ most likely tokens of the target model, similarly to how it is defined in Bachmann et al. (2025). We report $K{=}2,4,8,\dots,|V|$ for different speed-accuracy tradeoffs: increasing $K$ results in more accepted tokens but reduces accuracy.
@@
The results in Figure 3 demonstrate that AutoJudge decoding can achieve substantial speedups. Notably, our algorithm can accept over 40 tokens per target model forward pass in 8-shot evaluations for the 8B draft / 70B target model pair with a ${\leq}1\%$ change in accuracy. Varying the classifier threshold allows us to achieve even greater speedups at the cost of several percentage points drop in accuracy. The heuristic-based top-$K$ baseline also achieves some speedups, but at the cost of a significantly higher accuracy drawdown. We report setups in Appendix D.1.
Figure 4: Downstream Pass@1 and the average number of accepted tokens on LiveCodeBench for (left) Llama-3.2 1B draft / Llama-3.1 8B target and (right) Llama 3.1 8B draft / Llama 3.1 70B target. We use Instruct versions for both model pairs and report additional details in Section 4.2.
@@ 4.2 Programming with LiveCodeBench
Next, we test if the AutoJudge search algorithm is able to generalize between domains. For this purpose, we evaluate the same model family on the LiveCodeBench (Jain et al., 2024) programming benchmark. Here, we use the code_generation_lite∗*∗**huggingface.co/datasets/livecodebench/code_generation_lite dataset with the version tag release_v5. The dataset contains 880 programming tasks; we evaluate on all three subsets: easy, medium and hard. Since LiveCodeBench does not have a dedicated training split, we evaluate using out-of-fold predictions. Namely, we split the dataset randomly into 5 folds. For each fold, we evaluate using the classifier trained on the 4 remaining folds. We use the standard evaluation protocol: extracting the generated code and evaluating it using the benchmark’s built-in test suite. We otherwise follow the standard evaluation protocol for this benchmark and report Pass@1 in zero-shot setting.
Similarly to Section 4.1, we use the training data to find important tokens — this time, in terms of the resulting program correctness, measured as passing tests. Since the calibration dataset is smaller and further subdivided into folds, we only have $\approx$}27K mismatching tokens to train the classifier (with a slight ${\leq}0.5K$ variation depending on the active fold). Furthermore, we found that only ${\approx}5\%$ of the mismatching tokens were deemed to affect the output quality. We provide example token assignments in Figure 2 (right): notably, the tokens deemed important in that case would not appear in GSM8K in the same context. We otherwise follow the same training and evaluation protocol as before.
@@
The results in Figure 4 are similar to what we observed in Section 4.1: AutoJudge decoding can accept over 22 tokens per speculative decoding cycle at the cost of a ${\approx}2\%$ accuracy decrease. This results in approximately $2{\times}$ increase over traditional speculative decoding (Leviathan et al., 2023). The top-$K$ baseline can similarly achieve some increase in the number of accepted tokens, but AutoJudge decoding offers significantly better quality-speed tradeoffs across all configurations. We report additional configurations and threshold values in Appendix D.2, including the setup for AutoJudge decoding “out-of-domain”, i.e. using the classifier trained on GSM8K for LiveCodeBench evaluation. This out-of-domain configuration results in inferior performance, which aligns with our hypothesis that the important tokens depend on the problem type and evaluation criteria.
4.3 Inference Benchmarks with vLLM
In this section, we report the GPU inference speed of AutoJudge classifiers integrated into vLLM speculative decoding. For these benchmarks, we use the same GSM8K evaluation setup as in Section 4.1 and evaluate 8B/70B and 8B/405B model pairs. We integrate AutoJudge classifier into vLLM’s built-in lossless speculative decoding based on Leviathan et al. (2023). We compare AutoJudge decoding against lossless speculative decoding with draft window size set to 32: for each configuration, we report the absolute speed (tokens per second), as well as the relative speedup with respect to each baseline. We run 8B/70B benchmarks on 4 A100-SXM GPUs with NVSwitch and 8B/405B on 8 H100-SXM5 80GB GPUs with NVSwitch. We provide configuration details in Appendix E.
Our results, reported in Table 1 demonstrate that the advantage of our algorithm in terms of accepted tokens (Sections 4.1 & 4.2) translates proportionally to real-world tokens per second. Namely, if we allow ${\leq}2\%$ change in accuracy, our vLLM inference with AutoJudge classifier can achieve $108.5$ tokens per second for 8B/70B model pair, which translates into up to $2.21\times$ speedup compared to regular speculative decoding. For 8B/405B model pair we achieve $56.1$ tokens per second. In Appendix F we report additional benchmark configurations with similar performance gains. Also, we provide evaluations with individually tuned window sizes in Appendix  G as well as evidence of vLLM inference inconsistency in Appendix  I.
Limitations.
Our approach assumes that the downstream task has a way to determine whether or not two solutions are equivalent. While this is true for many tasks, it is less obvious for tasks such as story writing, image editing or general conversation. While it is possible to run AutoJudge on these tasks with an LLM judge to determine if the two answers are equivalent, we did not test this in our experiments. Additionally, running Algorithm 1 consumes compute resources, mostly LLM inference, in the form of local inference or API calls. We discuss this in more detail in Appendix J.
Table 1: Inference benchmarks (with vLLM) of AutoJudge classifiers on the GSM8K dataset for (left) Llama 3.1 8B draft / 3.1 70B target and (right) Llama 3.1 8B draft / 3.1 405B target models. Bolded columns highlight configurations that achieve strong speedups with minimal (1-2%) accuracy degradation.
Llama 3.1 8B draft / 3.1 70B target
Threshold 0.005 0.031 0.145 0.230
Accuracy, % 92.0 91.9 89.9 88.0
Speed, tokens/s 63.8 94.6 108.5 108.0
Speculative Decoding:   49.2 tokens/s
Speedup(ours) 1.30 1.92 2.21 2.20
Llama 3.1 8B draft / 3.1 405B target
Threshold 0.01 0.04 0.07 0.14
Accuracy, % 96.7 94.9 93.4 90.7
Speed, tokens/s 47.7 56.1 59.2 61.5
Speculative Decoding:   44.5 tokens/s
Speedup(ours) 1.07 1.26 1.33 1.38
5 Conclusion
@@
In this work, we propose and evaluate a fully automated protocol for task-specific acceleration of speculative decoding. Our initial experiments suggest that a simple procedure can successfully determine which of the mismatching tokens in the LLM response affect the downstream quality for both mathematical reasoning and programming tasks. We hope that AutoJudge can facilitate the use of Judge Decoding across different task types, languages and modalities.
@@
In future work, we aim to explore the performance of AutoJudge decoding across additional tasks and models and analyze how its outputs differ depending on the use case. Since our approach does not require human annotation, another research direction is to apply our algorithm to find important “tokens” for more complex openended tasks such as dialog or creative writing. While it is challenging to directly compare against the original Judge Decoding (see the end of Section 3), it would nonetheless be interesting to compare human and automated annotations in this setting. We also plan to explore how AutoJudge can be combined with more advanced speculative decoding algorithms, such as speculative decoding with tree-based drafts (Miao et al., 2023; Chen et al., 2024; Svirschevski et al., 2024) or learned drafting heads (Cai et al., 2024; Li et al., 2024b, 2025).
Acknowledgements
We would like to express our sincere gratitude to Denis Mazur for his valuable contributions to the implementation of API calls used in Algorithm 1 and for supporting inference with the Llama 405B model. We are also thankful for his positive influence on the overall atmosphere and team morale throughout the course of this project.
Appendix A Additional Considerations for Section 3.1
@@ Acknowledgements
$y_{next}=\underset{i}{\arg\max}\log P(i|x\oplus y_{1:t},\theta)+\textsc{GumbelPRNG}(s\oplus x\oplus y_{1:t}),$
where GumbelPRNG is a function that samples a pseudo-random variable from standard Gumbel distribution based on an input seed $s\oplus x\oplus y_{1:t}$. To recall, $\oplus$ denotes concatenation. This way, $y_{next}$ is distributed as $P(y_{next}|x\oplus y_{1:t},\theta)$, but it is deterministic when conditioned on the random state $s$. Hence, we sample a random state $s$ once at the beginning of Algorithm1, the entire procedure after that will also be conditionally deterministic (given $s$). We plan to test this in further versions of the paper.
@@ Appendix B Additional Details on Classifier Training
For the first set of experiments (Figure 5), we compare regularizer coefficients for Logistic Regression (left). We also report different classifier types: Logistic Regression, a Random Forest with 128 trees and a multi-layer perceptron (MLP) with a single hidden layer consisting of 128 hidden units with ReLU activation. For consistency, we run all models using Scikit-Learn [Pedregosa et al., 2011] v1.4.2 with all other settings kept to their default values. For LogisticRegression we set max_iter parameter to $500$. For MLP, we perform early stopping on yet another 10% subset of the training set with built-in default MLPClassifier early stopping parameters. For RandomForest we use min_samples_leaf set to $0.001$. For this evaluation, all classifiers use draft and target model hidden states (concatenated) encoding the draft token, which is our main setup from Section 3.2.
@@ Appendix C Precision Matters for Speculative Decoding
Appendix D Additional Evaluations
D.1 Additional Evaluations for Section 4.1
In Figure  8, we report accuracy and the number of accepted tokens for Llama-3.1-8B draft / Llama-3.1-70B target and Llama-3.1-8B draft / Llama-3.1-70B target model pairs in GSM8K 8-shot setup.
Figure 8: Downstream accuracy and the average number of accepted tokens for GSM8K 8-shot with Llama-3.1-70B-Instruct target and Llama-3.2-1B-Instruct draft models (left) and Llama-3.1-70B-Instruct target and Llama-3.1-8B-Instruct draft models (right)